OpenStudy (anonymous):

If the length of each side of a cube is tripled, is the surface area tripled? Use an example to support your answer

OpenStudy (anonymous):

the surface area would be 9 times larger

OpenStudy (anonymous):

as an example use a cube with side = 1. the surface area would be: SA = 1*1*6 = 6 the uses a side = 3. SA = 3*3*6 = 54 54/6 = 9